OpenStudy (anonymous):

Use algebra to find the inverse of the given one-to-one function. f(x) = (x^5 + 10)^3

OpenStudy (anonymous):

form your equation with x's and y's: \[y=(x^5+10)^3\] now replace your x's and y's: \[x = (y^5+10)^3\] and solve this altered equation for y. <--- that is your inverse equation

OpenStudy (anonymous):

right I got to there and the thing that is confusing me is the\[y^{5}\]

OpenStudy (anonymous):

Roots and Powers are inverse functions, you can clear power of 5 by taking the 5th root, for example: \[x=y^5\] \[\sqrt[5]{x}=\sqrt[5]{y^5}\] \[\sqrt[5]{x}=y\]